### v1

# MOSS: Motion-based 3D Clothed Human Synthesis from Monocular Video

Abstract
Single-view clothed human reconstruction holds a central position in virtual reality applications, especially in contexts involving intricate human motions. It presents notable challenges in achieving realistic clothing deformation. Current methodologies often overlook the influence of motion on surface deformation, resulting in surfaces lacking the constraints imposed by global motion. To overcome these limitations, we introduce an innovative framework, Motion-Based 3D Clothed Humans Synthesis (MOSS), which employs kinematic information to achieve motion-aware Gaussian split on the human surface. Our framework consists of two modules: Kinematic Gaussian Locating Splatting (KGAS) and Surface Deformation Detector (UID). KGAS incorporates matrix-Fisher distribution to propagate global motion across the body surface. The density and rotation factors of this distribution explicitly control the Gaussians, thereby enhancing the realism of the reconstructed surface. Additionally, to address local occlusions in single-view, based on KGAS, UID identifies significant surfaces, and geometric reconstruction is performed to compensate for these deformations. Experimental results demonstrate that MOSS achieves state-of-the-art visual quality in 3D clothed human synthesis from monocular videos. Notably, we improve the Human NeRF and the Gaussian Splatting by 33.94% and 16.75% in LPIPS* respectively. Codes are available at https://wanghongsheng01.github.io/MOSS/.
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding Author.
Keywords: 3D Gaussian Splatting, human reconstruction, matrix-Fisher.
Introduction
Figure 1: MOSS reconstructs 3D clothed humans with detailed joints and fine clothing folds. The right image demonstrates that MOSS surpasses the visual quality of previous works on MonoCap. (LPIPS* = LPIPS × 103). Larger circles denote higher FPS.
The significance of accurately modeling clothed humans is on the rise, especially in industries such as fashion, entertainment, and computer games [44, 16, 1]. When striving to create lifelike mannequins, one of the most significant challenges lies in accurately depicting the dynamic movements of the human body.
Recently, 3D Gaussian Splatting(3DGS)  [9, 3, 23, 38, 22, 4, 27] has greatly improved the efficiency and quality of 3D reconstruction by sampling object geometry through Gaussian distributions. Existing work on reconstructing the human body with 3DGS [15, 25] introduces the SMPL as the human body prior, which can restore a more realistic human body. However, these works neglect the hierarchical structure constraints of the kinematic tree [11] and the constraints of global motion information in reconstructed human surface, resulting in blurred joint details. Besides, the restored surface deformation is still underexplored. Another line of works [48, 49] developed normals to model clothing to characterize the deformation superficially. The influence of mutual occlusion between clothing folds on the fold rendering was not considered, resulting in the loss of clothing folds on the reconstructed human surface.
To confront the present difficulties, we introduce an innovative framework Motion-Based Clothed 3D Humans Synthesis (MOSS). MOSS commences by examining the origin of the surface deformation. We use the motion factors (displacement and rotation) from the kinematic tree for Gaussian control to improve the human body reconstruction in large-scale motion.
First, with the aim of deformation reconstruction, we propose KGAS, a module that extracts the human body from decomposing the matrix-Fisher distribution parameter, obtaining the principal axis concentration and rotation factor to explicitly control Gaussians for 3DGS rendering of human surface deformation. During the Gaussian layout process, the principal axis concentration performs as a density factor that corrects the sampling probability of Gaussian split to obtain surface deformation-aware Gaussians. In the following split control, the principal axis concentration and the rotation factor dynamically adjust the direction and scale of the Gaussian, realizing that each Gaussian expresses the clothed human body under the constraint of global motion information.
In single-view videos, local deformations often obscure each other, resulting in insufficient supervision of pixel features in 2D images. Consequently, the local deformations may fail to be reconstructed. To address this issue, it is imperative to consider mitigating the problem of local occlusion by capturing local geometrical deformations. We introduce the UID, where the core lies in the covariance matrix, which characterizes the heterogeneity of the spatial structure of the local point cloud. This characterization facilitates the discernment of local density and morphological differences within the point cloud. UID is grounded in explicit human structure retrieval from the surface, enabling localization to areas exhibiting drastic deformations. Subsequently, geometric reconstruction is performed to compensate for these deformations, thereby facilitating the restoration of occluded local geometry.
Based on the synergistic work of KGAS and UID, our method can improve the reconstruction quality of the clothed human in scenes with substantial human motion, while ensuring the efficiency of training and rendering. Our contributions are as follows:
(1) MOSS creatively extracts global motion factors from the origin of surface deformation. Employing these factors to guide Gaussians to the finest granularity restores realistic joint details and fine clothing folds in large-scale motion.
(2) We propose KGAS to propagate motion information obtained from the kinematic tree to all Gaussians, which effectively guides the 3D Gaussians, accurately restoring the surface deformation.
(3) We propose UID to detect variations of the orientation of local Gaussians, identifying large deformations to articulate the geometric deformation of the human body surface in a directional manner.
(4) Experimental results show that MOSS achieves state-of-the-art human visual perception performance with real-time rendering.
1 Related Work
1.1 3D Reconstruction and Rendering.
In the domain of 3D reconstruction and rendering, NeRF [31, 6, 21, 43, 34, 32, 51, 50, 30, 18, 12] has been recognized for its ability to enable high-quality rendering. However, NeRF often requires significant time and resources for computation. Recently, 3D Gaussian Splatting [9, 3, 23, 38, 22, 4, 27] has emerged as a promising method for point cloud rendering. By incorporating a spherical harmonic function to represent color attributes and facilitating efficient computation, 3DGS enhances the perception of geometric structures through pixel-level Gaussians. This advancement significantly enhances the efficiency and quality of 3D reconstruction processes.
1.2 Human Reconstruction.
In the field of human modeling, many methods have been used to model human shape and pose. One of the commonly used methods is SMPL [28], which is a vertex-based additive human body model that uses linear algebraic models to model shape and pose separately. At this stage, based on 3DGS, some methods [2, 5, 7, 29] use an explicit human structure SMPL priori to aid in the 3DGS optimization process to reconstruct diverse and realistic 3D human models with high-quality. PSAvatar [52] uses discrete geometric primitives to create parametric deformable shape models that are represented in fine detail and rendered with high fidelity with the help of 3D Gaussian. Human101 [25] uses human-centered forward in the SMPL prior to deforming 3D human models. Gaussian animation method to deform the parameters of 3D Gaussian and improve the rendering speed. GauHuman [15] not only introduces SMPL prior, but also further accelerates the 3DGS by introducing KL dispersion. These methods can restore a more realistic human appearance. However, they lack the constraints of the SMPL kinematic tree when considering human joint linkage effects and surface deformation, which can lead to problems such as blurred joint edges or confusion between joints and clothing when reconstructing complex human poses or when there is clothing occlusion.
1.3 Surface Reconstruction.
In the realm of reconstructing human clothing surfaces, traditional template mesh models have limitations in capturing clothing details, some approaches use neural representations [39, 40] to handle complex details like hairstyles and clothing, which considers the human body surface as an implicit function, while utilizing neural networks to learn parameters and effectively model the complex details. Other [48, 49] specifically utilize normals to handle extreme surface poses. Recent advances [31, 6, 21, 43, 34, 32] include the use of neural networks to represent dynamic human models and the extension of NeRF to dynamic scenes. Additionally, GaussianBody [26] proposes a segmentation plus scaling strategy to improve point cloud density, and solve the texture blur problem by a pose refinement approach. HDHumans [53], jointly optimize neural implicit fields and explicit template meshes, capable of handling loose clothing but with slower rendering speeds.
Figure 2: MOSS framework. MOSS rotates and zooms the Gaussians with Fisher. The T-pose is converted to the target pose and the surface folds are refined.
2 Method
Overview. In this paper, we aim to efficiently reconstruct high-quality clothed humans from single-view videos involving big movements. Our approach builds on the assumption that the corresponding SMPL parameters $\beta$, $\theta$, the calibrated camera parameters, and the human region mask are all given. Figure 2 illustrates our framework. We first use joint rotations to control the spatial position and orientation during the 3D Gaussian density transformation (Section 2.2, Section 2.3). The density of the Gaussian distribution is captured by the matrix-Fisher distribution, and the transformation of Gaussians is guided by density. Additionally, significant surface deformation is identified by analyzing the direction of maximum surface geometric variations (Section 2.4).). MOSS distinguishes between the human body and its surface, and constrains the newly generated human body surface Gaussian with whole-body motion information. This enables the reconstruction of more realistic and reasonable details of human surface deformation.
2.1 Preliminary
SMPL and LBS Transformation. The SMPL human model comprises a human skeleton with $N=6890$ vertices and $k=24$ joints. In this context, let’s define the joint rotation as $\theta_{joint}=[\theta_{joint_{0}}^{T},...,\theta_{joint_{K}}^{T}]^{T}$, where $K=23$, and $\theta_{joint_{k}}\in R^{3}$ denotes the rotation of part $k$ relative to its parent in the kinematic tree. Additionally, $\theta_{joint_{0}}$ refers to the orientation of the root part.
With Linear Blend Skinning(LBS), the position $p^{\prime}_{i}$ of the $i$th vertex $p_{i}$ in canonical space after skinning is denoted as: ${p}_{i}^{\prime}=\sum\nolimits_{k=1}^{K}w_{ki}G_{0k}(\theta_{joint},J(\beta)){p}_{i}+G_{1k}(\theta_{joint},J(\beta))$, where $G_{0k}(\theta_{joint},J(\beta))$ is the rotation transformation matrix of the joint $k$ that depends on the joint rotation $\theta_{joint}$ and the pose change parameter $\beta$ of human. Similarly, $G_{1k}(\theta_{joint},J(\beta))$ is the translation transformation matrix. $w_{ki}$ is the skinning weight of the joint $k$ on vertex $i$, which determines the extent to which joint $k$ affects vertex $i$.
Gaussian Splatting. We use $f_{\mathrm{3DGS}}=\sum\nolimits_{i}G(x,y,z,\mu_{i},\Sigma_{i})\cdot c_{i}$ to express 3DGS, where $c_{i}$ denotes color. where $G$ is a Gaussian function with inputs as a center (position) $x,y,z$, mean $\mu_{i}$ and covariance $\sum\nolimits_{i}$:
$G(x)=e^{-\frac{1}{2}x^{T}\Sigma^{-1}x},\Sigma=R_{g}S_{g}^{T}S_{g}R_{g}^{T}.$ (1)
Matrix-Fisher distribution. All 3D rotation matrices lie on a nonlinear and closed manifold SO(3). The matrix-Fisher distribution is defined over SO(3) [45, 20, 8]. For a given random rotation matrix $\textbf{R}\in SO(3)$, 3D special orthogonal groups can be defined as $SO(3)=\{\textbf{R}\in\mathbb{R}^{3\times 3}|=\textbf{R}^{T}\textbf{R}=\textbf{I},det(\textbf{R})=1\}$. The matrix-Fisher distribution is obtained according to the matrix rotation parameter denoted as $\theta_{joint_{i}}$. The probability density function of $R$ is given by the matrix-Fisher distribution:
$p(\textbf{R}|\theta_{joint_{i}})=\frac{1}{c(\theta_{joint_{i}})}\exp\left(\text{tr}\left(\theta_{joint_{i}}^{T}\textbf{R}\right)\right)=\mathcal{M}(\textbf{R};\theta_{joint_{i}}).$ (2)
In this equation, $c(\theta_{joint_{i}})$ is a normalization constant that ensures that the total probability sums to 1 on $SO(3)$. We use matrix-Fisher to characterize the joint rotation probability distribution.
The derivation reveals that the most probable rotational state $R_{mode}$ of a given joint (see the Appendix for a detailed derivation) can also be represented by the elements obtained by decomposing $\theta_{joint_{i}}$ by the proper singular value:
$R_{mode}={\text{arg max}}\ p\left(R|\theta_{joint_{i}}\right)=\textbf{U}\textbf{V}^{T}with\ R\in\text{SO}(3).$ (3)
2.2 Global Motion Extraction
To restore the joint details of the clothed human body, one of the challenges is to incorporate the SMPL kinematic tree. The SMPL kinematic tree constrains the joint changes of the human body. To this end, we guide the 3D Gaussians with global motion, the source of surface deformation, from SMPL kinematic tree, which constraints the new Gaussians precisely and to the finest granularity (Section 2.3). In this section, our first use of principal axis concentration matrices linked to scale transformations of Gaussian distributions is a breakthrough in the exploration of the semantics of Fisher distributions.
Matrix-Fisher over the SMPL kinematic tree. Implementing 3DGS for clothed humans, the Gaussians are typically weakly constrained by the entire body. We first aim to extract joint motion information. Inspired by  [24, 33, 41], we attempt to utilize the joint rotation distribution to extract global motion and subsequently employ it to guide the 3D Gaussian density transformation.
We perform a proper singular value decomposition $\mathbf{\theta_{joint}}$ of the matrix parameters $\mathbf{\theta_{joint}}$ of the joint rotation distribution $\mathbf{\theta_{joint}}=\boldsymbol{U}\boldsymbol{S}\boldsymbol{V}^{T}$. where $\boldsymbol{U}$ and $\boldsymbol{V}$ are orthogonal matrices, $U$ characterizes the principal axes, and $S=diag(s_{1},s_{2},s_{3})$ is a diagonal matrix consisting of the singular values $s_{1},s_{2},s_{3}$ of $\mathbf{\theta_{joint}}$ $s_{1}>s_{2}>s_{3}$ as a diagonal matrix. The singular values can dominate the degree of concentration of the distribution along the principal axes. It is considered a scaling factor that affects the scaling matrix $S_{g}$ of the Gaussian distribution.
The most probable rotation $R_{mode}$ of the joints is obtained by matrix-Fisher’s mode (Section 2.1), which contains joint motion information. To make new Gaussians correspond with the SMPL kinematic tree’s comovement, we use $R_{mode}$ to provide direction guidance for the rotation $R_{g}$ of the new 3D Gaussians.
Thus, we complete the human joint motion information extraction, obtaining the two factors $S$ and $R_{mode}$ to be used in Section 2.3.
Joint-Driven Orientation Refinement. After defining the matrix-Fisher over the SMPL kinematic tree, we further extract the global motion information during human pose transformation and utilize it to guide the displacement of Gaussians. We propose Joint-Driven Orientation Refinement(JNTM), which learns the correlation weights of joints on the influence of all Gaussians, and explicitly controls the Gaussians to realize ”pulling one to move the whole body”. We use Joint Driver in JNTM to propagate motion information to every Gaussian. Technically, we implement Joint Driver via cross-attention for the sake of simplicity and directness.
To characterize the joint linkage information of the whole body, we introduce the grandfather node relationship of the kinematic tree to constrain the matrix-Fisher distribution of each joint. We model the matrix-Fisher distribution of the rotation matrices of the child nodes of a kinematic tree after the rotation matrices of its parent nodes are known: $\mathcal{M}\left(\theta_{joint_{i}};R_{i}\right)=p(R_{i}|\theta_{joint_{j}},j\in p(i)),$ where $\theta_{joint_{i}}$ denotes the $i$th joint rotation parameter to be estimated. $\theta_{joint_{i}}$ can be estimated more exactly by combining the kinematic tree and introducing the prior information of the parent nodes in $p(i)$ by autoregression. The estimated child node $\theta_{joint_{i}}=f_{i}(\theta_{joint_{i}},\{\theta_{joint_{j}}\}_{j\in P(i)})$.
The matrix-Fisher over the SMPL kinematic tree dynamically represents the real-time global motion which serves as guidance for new Gaussians. The parameters of the matrix-Fisher distribution are estimated by autoregression.
2.3 Motion Guidance in 3D Gaussian Density Control
Figure 3: Fisher’s Gaussian Sampling. This is a 2D example of how spindle concentration affects Gaussian sampling. note that the color bar represents the probability of sampling, with darker colors representing higher probabilities.
When performing large-scale motion, the relative displacement of limbs is usually larger than that of the trunk. However, Gaussians learning the limb-aroused cloth folds around the trunk are produced by human trunk Gaussians, lacking essential global motion information. Hence, we leverage global motion extracted from Section 2.2 to guide the density transformation process of Gaussians for better capture of the human body under substantial motion.
Density Perceptual Sampler. The scaling factor of 3DGS is corrected with the principal axis concentration matrix $S$ to change the size of the sampling area in order to make it closer to the real surface shape. This contributes to a more accurate representation of the surface deformation by the Gaussians. The corresponding Gaussian distribution obtained after the adjustment by $S$ is $g(x)=e^{-\frac{1}{2}x^{T}\Sigma^{\prime-1}x}$, as shown in Figure 3. The displacement of the new Gaussian denoted as $\Delta\textbf{x}$ is generated from $g(x)$, and the position of the center of the new Gaussian ellipsoid $\textbf{x}_{new}$ is obtained from the original Gaussian by this displacement transformation: $\textbf{x}_{new}=\Delta\textbf{x}+\textbf{x}_{source},$ so the new Gaussian obtained by cloning from the original is $G_{new}=G(x_{new})=e^{-\frac{1}{2}x_{new}^{T}\Sigma^{-1}x_{new}}.$ The locations of the resulting new Gaussians are closer to the surface, allowing these Gaussians to more realistically represent the geometric deformation of the surface.
Motion Guidance for Gaussians. Next, we use the $S$ and $R_{mode}$, which contain global motion information, to control the density transformation process of the Gaussians. The rotation probability concentration matrix $S$ characterizing the different principal axes is used as a scale correction factor for the Gaussian to control the size of the Gaussians. The joint rotation state estimation $R_{mode}$ is used to control the direction of rotation of the Gaussians. matrix-Fisher extracts real-time motion information that is used to guide the Gaussians to learn human posture changes in real-time. In the Gaussian clone and split process, the estimated $S$ and $R_{mode}$ guide the scaling and rotation of the new Gaussians obtained after cloning, respectively: $S_{g}^{\prime}=mat\_multi(S,S_{g})$, $R_{g}^{\prime}=mat\_multi(R_{mode},R_{g}).$
$g(x_{new})=e^{-\frac{1}{2}x_{new}^{T}\Sigma^{\prime-1}x_{new}}$ and $\Sigma^{\prime}=R_{g}S_{g}^{\prime T}S^{\prime}_{g}R_{g}^{T}$ are obtained, corresponding to the Gaussian distribution. Consequently, KGAS (Section 2.2, Section 2.3) achieves Gaussians density perception. Besides, the new Gaussians can be effectively constrained by the whole-body motion, leading to the generation of more realistic and plausible human body and surface details.
2.4 Surface Deformation Detector
Figure 4: Solving occlusion problems with UID (2D). There is a potential problem of smaller folds being occluded by obvious folds due to the viewing angle. By calculating the degree of directional change in the local distribution of Gaussians, the regions with large deformation on the surface are localized and densely processed.
Small folds contribute little to the total pixel-wise loss, leading to weakened supervision. In addition, the mutual occlusion among folds even toughens supervision in single-view. Although  [19, 15, 25] pick out Gaussians far from the real human with positional gradients, these are not all for joint details or cloth folds. For this reason, we propose the Gaussian clustering-based UID to detect complex or large surface deformations, see Figure 4.
We densify the Gaussians based on gradient localization to increase the rendering granularity for better reproduction of clothing folds and joint details. Remember that the set of Gaussians selected by the gradient is $M=\{m1,m2,...\}$. For each point $m_{i}$ in the point set $M$, select the $k$ Gaussians that are closest to it to form the point set $M_{i_{K}}$. The local center of mass C of $M_{i_{K}}$ is given by the following equation:
$C=\left(\frac{1}{n}\sum\nolimits_{i=1}^{n}x_{i},\frac{1}{n}\sum\nolimits_{i=1}^{n}y_{i},\frac{1}{n}\sum\nolimits_{i=1}^{n}z_{i}\right),$ (4)
where $n$ is the total number of points in the point set $M_{i_{K}}$ and $(x_{i},y_{i},z_{i})$ denotes the coordinates of each point.
We compute the covariance matrix of these points in $m_{i}$ with its neighborhood. Assuming that the local center of mass is $\bar{m}$ and the points in the neighborhood are $m_{i_{k}}$, the covariance matrix Cov of the sample is given by the following equation:
$\operatorname{Cov}=\frac{1}{k-1}\sum\nolimits_{i=1}^{k}\left(m_{i_{k}}-\bar{m}\right)\left(m_{i_{k}}-\bar{m}\right)^{T}.$ (5)
Eigen-decomposition of the covariance matrix is performed to obtain the eigenvalues and eigenvectors:
$\operatorname{Cov}\cdot v_{i}=\lambda_{i}\cdot v_{i},$ (6)
where $v_{i}$ is the eigenvector and $\lambda_{i}$ is the corresponding eigenvalue. The eigenvector corresponding to the smallest eigenvalue of the covariance matrix, $\mathcal{N}_{i}$, points in the direction of the smallest change in the data  [46]. To accurately capture the location of the large deformation, we use the $\mathcal{N}_{i}$ for each selected Gaussian to calculate the angle $\theta\left(\mathcal{N}_{i},\mathcal{N}_{j}\right)$, which characterizes the magnitude of the localized geometric deformation of the surface. Specifically, it is computed as follows, given the $\mathcal{N}_{i}$ of each Gaussian $m_{i}$ in the set of points with its neighborhood $N_{i}$ of the other points $m_{j}$ in the set of points $\mathcal{N}_{j}$:
$\theta\left(\mathcal{N}_{i},\mathcal{N}_{j}\right)=arccos\frac{<\mathcal{N}_{i},\ \mathcal{N}_{j}>}{||\mathcal{N}_{i}||||\mathcal{N}_{j}||},j\in N_{i}.$ (7)
If $\theta\left(\mathcal{N}_{i},\mathcal{N}_{j}\right)$ is greater than a threshold, it is considered as a location with large surface deformation, which is labeled and used to guide the generation of more Gaussians at that location to fully learn the human surface deformation. We precisely capture the hard-to-render surface deformations with folds or protrusions by the minimum direction of geometric deformation and densify them to better reproduce clothing folds and joint details.
2.5 Structure Perception Optimization
Existing methods use random or Structure-from-Motion (SfM)  [42] to initialize 3D Gaussians. We directly sample 6890 Gaussians from the SMPL for initialization to reduce the cost of training. During inference, LBS weights and pose parameters saved in the training phase are used to improve speed while maintaining high rendering quality.
Loss. Combining image generation quality, human perception, and joint rotation constraints, we adopt the following Loss:
$L=\lambda_{1}L_{image}+\lambda_{2}L_{percep}+\lambda_{3}L_{joint}.$ (8)
Among them, $L_{image}$ is the loss used to supervise the quality of image generation, integrating color, and mask. $L_{percep}$ is for human vision perception. $L_{joint}$ constrains the estimation of joint rotational, which enables more accurate parameter estimation for matrix-Fisher distribution (see appendix for specific realizations).
2.5.1 Implementation Details.
We employ AdamW with different parameters for different modules. Experiments are conducted on a single RTX 3090 GPU. The model takes about 3,000 iterations or so and finally converges within 4 minutes (512 $\times$ 512 resolution). Furthermore, a comprehensive overview of our network structure and hyperparameters can be found in the appendix.
3 Experiments
3.1 Experimental Setups
Datasets.
We conduct our experiments on ZJU-MoCap [36] and MonoCap datasets. ZJU-MoCap includes humans performing complex motions such as licking, punching, etc. It contains foreground masks, camera and SMPL parameters. We pick one camera for training and others for testing. MonoCap dataset is processed by  [37] from DeepCap [13] and DynaCap [14] dataset. It also provides camera and SMPL parameters as well as human masks. We use one camera view for training and select ten uniformly distributed cameras for testing.
Comparison methods.
Compare MOSS with two categories of SOTA methods: Human NeRF-based methods, such as NeuralBody [36], HumanNeRF [47] AnimateNeRF [35], InstantNVR [10] InstantAvatar [17]. 3DGS methods, including GauHuman [15] and Human101 [25]. Following  [10], for a fair comparison, we do not optimize the test time for the SMPL parameters using images from the test set on InstantAvatar [17]. More details see the appendix.
Metrics. PSNR, SSIM, and LPIPS* are utilized as indicators of quality, while FPS is employed as a metric for speed. PSNR evaluates the variance between pixel values of two images. SSIM measures image similarity in terms of luminance, contrast, and structural integrity. LPIPS* compares two images in the feature space, which is considered to be closer to human visual perception.
3.2 Evaluation
To validate the effectiveness of MOSS in solving the problem of lacking global constraints for clothed human reconstruction, we compare our method with some previous human reconstruction methods [36, 47, 35, 10, 17, 15, 25] on the ZJU-MoCap and MonoCap datasets. Our method outperforms the comparison methods on most of the metrics, as shown in Table 1, thus validating the effectiveness of MOSS in solving the problem. Detailedly, MOSS achieves state-of-the-art LPIPS* on ZJU-MoCap and MonoCap, and the PSNR also achieves state-of-the-art on the Monocap dataset. It quantitatively demonstrates that our approach can maintain excellent rendering results while achieving industry-leading FPS.
GauHuman initializes with SMPL and prunes the 3D Gaussians with large KL divergence in adaptive density control, which only provides static human body information. On the contrary, we propose KGAS to control the Gaussian Splatting through the density and rotation factors. Compared with GauHuman, our method improves the LPIPS* by 16.75% and 12.51% on the ZJU-MoCap and MonoCap datasets, respectively, and improves the PSNR by 0.11 on the MonoCap dataset, which illustrates the effectiveness of our method.
Compared with InstantNVR, our method improves the LPIPS* on ZJU-MoCap and MonoCap datasets by 33.94% and 29.98%, respectively. The PSNR on the MonoCap dataset improves by 0.95. InstantNVR manually divides the human body into multiple parts based on SMPL, mapping the vertices to the UV coordinate and the in-depth learning of each part. Although this effectively enhances the structural information of the image, it still lacks objectivity due to the manual division and the human body cannot be divided finely as well. On the contrary, the UID we designed based on KGAS can detect the local position with large deformation by objectively searching the surface and strengthening the supervision.
Reconstruction Efficacy of MOSS. We validate the constraint effect of MOSS on 3D Gaussians, as well as its ability to identify local regions with significant surface deformations. As shown in Figure 5, we select the same human body from two different viewpoints and the qualitative comparisons. The details of the human body joints are marked by the red box, and the details of the clothing folds and other deformations are marked by the yellow box. It shows that our method can reproduce the surface details and clothing folds with high quality in large motion scenarios.
Table 1: Quantitative evaluation of our method compared with previous works. We conduct experiments on the ZJU-MoCap and MonoCap datasets and list the average of metrics at a resolution of 512 $\times$ 512.
Method ZJU-MoCap MonoCap
PSNR↑ SSIM ↑ LPIPS*↓ FPS PSNR↑ SSIM ↑ LPIPS*↓ FPS
NeuralBody 29.03 0.964 42.47 1.48 32.36 0.986 16.70 0.98
HumanNeRF 30.66 0.969 33.38 0.30 32.68 0.987 15.52 0.08
AnimateNeRF 29.77 0.965 46.89 1.11 31.07 0.985 19.47 0.31
InstantNVR 31.01 0.971 38.45 2.20 32.61 0.988 16.68 0.75
InstantAvatar 29.73 0.938 68.41 4.15 30.79 0.964 39.75 2.54
Human101 31.79 0.965 35.75 104 32.63 0.982 16.51 104
GauHuman 31.34 0.965 30.51 189 33.45 0.985 13.35 189
MOSS (Ours) 31.01 0.963 25.40 276 33.56 0.985 11.68 257
Figure 5: To ensure a fair comparison, we compare  [36, 47, 35, 10, 15, 25] at 512 $\times$ 512 resolution. Our model shows better visual quality and more detail. More comparative results are detailed in the appendix.
Gaussian Guidance. As shown in Figure 5, MOSS achieves more excellent visual quality in the details of human joints. The edges of clothing and joints are clearly demarcated, and the orientation of clothing folds is closer to the real occasion (e.g., shorts and thighs). Among the compared methods, NeuralBody encodes human body structure with the latent code of SMPL vertices, but it cannot capture complex human movement. This results in blurred fists and incomplete hands in the figure. AnimateNeRF expects to overcome the ill-posed problem. The human joint skeleton is introduced as regularisation to facilitate training. Its rendering results still have artifacts in arms and calves. GauHuman’s Gaussians lack global constraints, leading to problem such as confusion between shorts and knees, as well as loss of hand details. The KGAS module of MOSS makes the human body reconstruction constrained by the topology of the kinematic tree and guides the 3D Gaussian displacements with the global motion, which is able to fully learn the motion of human, and the differences between the clothes and the human body. In addition, the orientation of clothing folds is also more realistic due to the constraints of motion direction.
Surface Deformation Detection. MOSS restores more surface details. On the contrary, InstantNVR’s modeling only considers the human body and ignores the clothing, resulting in a complete loss of clothing folds of the human reconstructed by InstantNVR. What’s more, the InstantNVR rendered shorts tightly clinging to the legs, which is different from the loose shorts in reality. The pixel-based loss used in GauHuman only provides weak supervision, which results in its restoration of only a few folds (e.g., the sides of the waistcoat and shorts on the torso shown in the yellow box in the first row of results). The UID detects complex deformations on the local body surface and compensates for the geometric reconstruction. Hence, UID effectively solves the problem of insufficient supervision in a single-view, and the folds that are hidden from each other can also be restored.
Table 2: We use Sequence 377 from ZJU-MoCap to test the effectiveness of Joint-Driven Orientation Refinement and its components.
Method LPIPS*↓ PSNR↑ SSIM ↑
Baseline 37.62 28.91 0.963
Hierarchical (Baseline + Autoregression) 32.55 30.12 0.971
JNTM (Hierarchical + Joint Driver) 24.98 30.87 0.971
Ours 15.18 31.84 0.974
Table 3: These evaluations are conducted using Sequence 377 from the ZJU-MoCap dataset, ensuring that each component of MOSS works.
Method LPIPS*↓ PSNR↑ SSIM ↑
Ours 15.18 31.84 0.974
w/o Density Perceptual Sampler 18.53 31.31 0.97
w/o Motion Guidance 16.89 31.75 0.973
w/o UID 20.71 31.07 0.968
3.3 Ablation Study
We conduct ablation studies to examine the role that every component plays in MOSS. The results of our experiments on the Sequence 377 from ZJU-MoCap are shown in Table 2, Table 3, Figure 6 and Figure 7. Table 2 presents the effectiveness of JNTM and its sub-components. The baseline is 3D Gaussian Splatting initialized with SMPL. By adding Autoregression and Joint Driver incrementally, the quality of human body reconstruction is progressively improved. Table 3 shows the results of MOSS and the outcomes after removing only one component each time. Through selectively disabling each component, we evaluate their contribution to the system. We will analyze the efficacy of individual components according to experimental results specifically in the following part.
Figure 6: Ablation studies result of rendered images. (a) Results of the JNTM are demonstrated. (b) Reflecting the validity by removing our proposed components.
Figure 7: Ablation studies result of depth images. Warmer tones indicate denser information and cooler tones indicate sparser information.
Joint-Driven Orientation Refinement. As is shown in Table 2, the Autoregression improves the performance in LPIPS*, PSNR, and SSIM. This validates that it is effective to use the kinematic tree to constrain the parameter estimation for the distribution of joint rotations. Furthermore, Joint Driver also improves the metrics, this is because it precisely controls Gaussians by joint movement. In comparison, Ours has significantly enhanced the LPIPS*, owing to the incorporation of guidance on Gaussians and surface deformation detection. The rendered and depth images in Figure 6 (a) and Figure 7 also show the progressively improved effect when sequentially adding subcomponents of JNTM.
Density Perceptual Sampler. We conduct the following ablation experiments to explore the Gaussian density and surface structure perceptual ability. The results are shown in Table 3 w/o Density Perceptual Sampler. According to the green box in Figure 6 (b), the absence of this component results in the blurred rendering of surface edges, especially the fingers and the face. It indicates that the Gaussians of the human body and surface are severely underdense on the boundaries in Figure 7. In comparison, the Density Perceptual Sampler uses the principal axis concentration to correct the sampling probability of the Gaussian split process, such that the sampling area is more aligned with the actual surface. Besides, the restored deformation is more significant and clear even if the human is performing motions such as head rotation, finger clenching, lumbar rotation, and foot padding.
Motion Guidance. The experimental results highlight the impact of Motion Guidance. Human body movements, such as fist clenching or arm raising, lead to local shadows. Comparison of blue boxes in Figure 6 (b) and Figure 7 shows that Motion Guidance can mitigate the confusion between shadows and actual body parts. Additionally, the blue circle illustrates that Motion Guidance helps to restore the precise shapes of irregular and flexible clothing (e.g., belt loop). This component guides Gaussians with global motion, dynamically adjusting the orientation and scaling of each Gaussian. Consequently, it not only enhances the fidelity of reproducing local shadows in human body motion but also improves the accuracy in describing texture and shape alterations in clothing. The efficacy of w/o Motion Guidance is further elucidated through the results of ablation experiments in Table 3.
UID. To verify UID’s ability to identify the complex folds on the surface, we perform the following ablation experiments, with the results shown in Figure 6 (b). If we omit UID and only use pixel-wise losses for supervision, it will lead to insufficient guidance. Especially in Figure 7, most of the body parts lack depth information. Hence, it gets more difficult to accurately reconstruct various human body surfaces (especially under occlusion scenarios, seen in Figure 4). Furthermore, a comparison between the metrics of w/o UID and Full in Table 3 reveals a decline in the performance across all evaluation metrics. Confusion between the human body and trunk clothing (see yellow boxes), surface rendering distortions (see the yellow circle), and significant loss of facial features, including drastic texture changes and loss of information in high-frequency regions (see the yellow polygon), are evident. In addition, Figure 6 also shows that without UID leads to inaccurate restoration of overlapping lower legs, knees, and shorts with significant artifacts.
4 Conclusion
To address the lack of global constraints for detailed reconstruction of clothed humans in motion, we propose MOSS for reconstructing humans from 3DGS guided by global motion. The framework sends the human body in motion prior to the Gaussian rendering process of the body surface and focuses on the locations where the surface deformation is significant. In our future work, we consider incorporating the graph to topologically guide 3D clothed human reconstruction. There is also a wide range of substantial human motion scenarios in many fields, such as virtual reality and the fashion industry, where our technology has potential applications. For instance, it could reduce game production costs, enhance player experience, and support fashion designers in optimizing their designs.